Problem 13-17

The U.S. Airfilter company has hired you as a supply chain consultant. The company makes air filters for residential heating and air-conditioning systems. These filters are made in a single plant located in Louisville, Kentucky, in the United States. They are distributed to retailers through wholesale centers in 100 locations in the United States, Canada, and Europe. You have collected the following data relating to the value of inventory in the U.S. Airfilter supply chain:

QUARTER 1
(JANUARY
THROUGH
MARCH)
QUARTER 2
(APRIL THROUGH
JUNE)
QUARTER 3
(JULY
THROUGH
SEPTEMBER)
QUARTER 4
(OCTOBER
THROUGH
DECEMBER)
All amounts in millions of U.S. dollars
Sales (total quarter):
United states 456 402 476 512
Canada 130 149 121 165
Europe 32 40 24 16
Cost of goods sold (total quarter) 377 362 384 431
Raw materials at the Louisville plant
  (end-of-quarter)
83 41 62 86
Work-in-process and finished goods at
  the Louisville plant (end-of-quarter)
197 141 140 226
Distribution center inventory
  (end-of-quarter):
United states 67 51 30 40
Canada 16 14 17 18
Europe 7 6 9 5


a. What is the average inventory turnover for the firm? (Round your answer to 3 decimal places.)

Average inventory turnover            


b. If you were given the assignment to increase inventory turnover, what would you focus on?

    

o Increase the amount of inventory or increase sales or both
o Reduce the amount of inventory or increase sales or both
o Reduce the amount of inventory or reduce sales or both


c. The company reported that it used $500M worth of raw material during the year. On average, how many weeks of supply of raw material are on hand at the factory? (Round your answer to 2 decimal places.)

Weeks of raw material supply [ ]
Answer #1

a)

Total COGS across all quarters = 377+362+384+431 = 1554


Raw Materials inventory = 83+41+62+86 = 272

Work-in-progress and Finished goods = 197+141+140+226 = 704


Distribution center inventory = (67+16+7) + (51+14+6) + (30+17+9) + (40+18+5) = 280

Average aggregate inventory = Raw Materials + Work-in-progress and Finished goods + Distribution center inventory

= 272+704+280
=1256

Average inventory turnover = COGS/Average aggregate inventory = 1554/1256 = 1.237261146 = 1.237 (Rounded to 3 decimal places)

b)

Correct answer is Reduce the amount of inventory or increase sales or both

Focus would be on : Reducing the WIP or distribution inventory that would have direct implication on reduction of average aggregate inventory .This will increase inventory turnover ratio.

Increase of sales also can increase inventory turnover ratio as it will increase COGS .

Decreasing inventory and increasing sales simultaneously also can increase inventory turnover ratio

c)

Weeks of supply of raw material = (Raw material inventory/COGS)*52 = ((272+500)/1554)*52 = 25.83269 = 25.83 (Rounded to 2 decimal places)